Listed here are the total costs associated with the production of 1,000 drum sets manufactured by TrueBeat. The drum sets sell for $540 each.   

Costs

Plastic for casing—$15,000

Wages of assembly workers—$90,000

Property taxes on factory—$8,000

Accounting staff salaries—$42,000

Drum stands (1,000 stands purchased)—$38,000

Rent cost of equipment for sales staff—$36,000

Upper management salaries—$205,000

Annual flat fee for factory maintenance service—$20,000

Sales commissions—$25 per unit

Machinery depreciation, straight-line—$35,000

  1. New Plant? Because TrueBeat’s capacity is limited to 8,000 units in their current production facility, management would like to explore the impact of building a new production facility. The company is considering a new production facility and equipment that will decrease their variable expenses by 30% but increase fixed expenses by 50%. TrueBeat still plans to produce and sell the same number of units in Year 4 (base). The new plant will give them a relevant range of 5,000 to 12,000 units.

  1. If the new production facility is built, what would be the company’s new (i) contribution margin per unit, (ii) fixed expenses and (iii) the new profit formula? Complete the table below.

  1. New CM per unit                __________________
  2. New Fixed expenses            __________________
  3. New profit formula             __________________

Under New production facility

New Break-even

New Year 4

Units

                           

6,000

Sales Revenue

NOI

-0-

  1. Assume TrueBeat sells the same units as planned for Year 4 in the new plant, calculate the new margin of safety in (i) units) (ii) dollars and (iii) percent.

  1. MoS units              ______________________
  2. MoS dollars            ______________________
  3. MoS percent   

  1. Conclusion: If you were a member of top management, would you have been in favor of constructing the new plant? Explain in 30 to 50 words
Answer #1

a. i. Current variable cost per unit = $ 168

New variable cost per unit = $ 168 x 0.70 = $ 117.60

New CM per unit = $ 540 - $ 117.60 = $ 422.40

ii. New fixed expenses = $ 346,000 x 1.50 = $ 519,000

iii. New profit formula :

NOI = 422.40 q - 519,000

Under new production facility
New Break-even New Year 4
Units 1,229 6,000
Sales revenue $ 663,660 $ 3,240,000
NOI $ 0 $ 2,015,400

New break-even in units = $ 519,000 / $ 422.40 = 1,228.69 units or 1,229 units

New break-even sales revenue = $ 519,000 / $ 422.40 * $ 540 = $ 663,660

NOI at sales level of 6,000 units = $ 422.40 x 6,000 - $ 519,000 = $ 2,015,400

MoS units = Actual Sales Units - Break-even Sales Units = 6,000 - 1,229 = 4,771 units

MoS dollars = ( 6,000 - 1,229 ) x $ 540 = $ 2,576,340

MoS percent = ( Actual Sales - Break-even Sales ) / Actual Sales = 4,771 / 6,000 = 79.52 %